0 0.1 1 0.1 2 0.5 3 4 0.1 The probability model below describes the number...

0 0.1
1 0.1
2 0.5
3
4 0.1

The probability model below describes the number of thunderstorms that a certain town may experience during the month of August. What should the value be in the missing cell?
Answer #1

we know thattotal probabilty should be equals to 1

P(0)+P(1)+P(2)+P(3)+P(4)=1

0.1+0.1+0.5+x+0.1=1

0.8+x=1

x=0.2